40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 10 14:57:01 2019, Koji, Summary, CDS, Channel recipe for iscaux upgrade D990694-B.pdfD990543A1.PDF
    Reply  Thu Jul 11 12:42:35 2019, gautam, Summary, CDS, P2 interface board IMG_7728.JPG
       Reply  Thu Jul 11 13:08:36 2019, Chub, Summary, CDS, P2 interface board 
          Reply  Thu Jul 11 13:09:22 2019, gautam , Summary, CDS, P2 interface board 
       Reply  Sat Jul 20 11:57:39 2019, gautam, Summary, CDS, P2 interface board IMG_7773.JPG
    Reply  Thu Jul 18 00:51:52 2019, Koji, Summary, CDS, iscaux electronics modifications LSC_whitening2.jpgLSC_whitening.jpg
       Reply  Thu Jul 18 22:34:40 2019, Koji, Summary, CDS, iscaux electronics modifications 
Message ID: 14770     Entry time: Thu Jul 18 00:51:52 2019     In reply to: 14744     Reply to this: 14775
Author: Koji 
Type: Summary 
Category: CDS 
Subject: iscaux electronics modifications 

Along with the plan in ELOG 14744, the ISC PD interface and the whitening filter board have been modificed. The ISC PD I/Fs were restored to the crate and the cables were connected. The whitening filteres are still on the electronics bench for some more tests before being returned to the crate.

The updated schematics were uploaded as https://dcc.ligo.org/D1900318 and https://dcc.ligo.org/D1900319

- Modification of the ISC PD interface: Jumpers between DIN96 P1 and P2. Replace all AD797s with OP27. In fact only I/F #1 (the left most)  had total 12 AD797 but the other units already had OP27s.

- Modification of the whitening filter: Jumpers between DIN96 P1 and P2.

Attachment 1: LSC_whitening2.jpg  7.451 MB  Uploaded Thu Jul 18 02:35:33 2019  | Show | Hide all | Show all
Attachment 2: LSC_whitening.jpg  7.756 MB  Uploaded Thu Jul 18 02:35:33 2019  | Hide | Hide all | Show all
LSC_whitening.jpg
ELOG V3.1.3-